OpenStudy (anonymous):

Zack ran t times around the track at school, which is 1/4 mile. He ran a total of 11/12 of mile. Solve for t.

OpenStudy (ranga):

If Zack ran around the track one time he would have run 1/4 mile. If he ran around the track "t" times, he would have run 1/4 * t miles. 1/4 * t = 11/12 solve for t.

OpenStudy (timaashorty):

What we know: 1time around the track = 1/4 mile So \[t \times\frac{ 1 }{ 4 }=\frac{ 11 }{ 12 }\] \[t=\frac{ 11 }{ 12 } \times \frac{ 1 }{ 4 }\] \[t = \frac{ 44 }{ 12 }\] \[t=\frac{ 11 }{ 3 }= 3 \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

:o

OpenStudy (timaashorty):

therefore; \[ \large t \ equals\ 3 \frac{ 2 }{ 3}\]

OpenStudy (anonymous):

is it possible to give 2 medals?

OpenStudy (anonymous):

Thanks!

OpenStudy (ranga):

On the second line, it should be t = 11/12 * 4 (not 11/12 * 1/4). But the final answer is correct. You are welcome. Give the medal to tima.

OpenStudy (timaashorty):

You're right my mistake. and welcome [=